OpenStudy (anonymous):

f(x)=|x+3|. Find all solutions to f(x-1)=x+4

OpenStudy (anonymous):

\[f(x)=|x+3|,f(x-1)=?\]

OpenStudy (anonymous):

the question on the homework is as stated. Where f(x-1)=x+4. I'm not sure where to start

OpenStudy (anonymous):

is it |x+3|-1=x+4 ?

OpenStudy (anonymous):

bno

OpenStudy (anonymous):

it is \[f(x-1)+|x-1+3|=|x+2|\]

geerky42 (geerky42):

\[f(\color{red}{x}) = |\color{red}{x}+3|\]So \(f(\color{red}{x-1})= |\color{red}{x-1}+3|=\underline{\underline{\underline{|x+2|}}}\)

OpenStudy (anonymous):

\underline\underline\underline i like it

OpenStudy (anonymous):

oh... so where does the x+4 come in?

OpenStudy (anonymous):

oh.. |x+2|=x+4 solve x

geerky42 (geerky42):

Yeah, so you "split" it into two cases, one where x+2 is positive and another one where x+2 is negative, so you have two equation to solve: \(x+2 = x+4\\x+2=-(x+4)\) Does that make sense?

OpenStudy (anonymous):

click. thank you. makes sense.